Peter is twice as old as Mary. Seven years ago the sum of their ages was 31. Find their ages.

Answer by Alan3354(69443)

Peter is twice as old as Mary. Seven years ago the sum of their ages was 31. Find their ages.
----------------
The sum of their ages now is 31 + 7 + 7 = 45
P + M = 45
2M + M = 45
--> 30 and 15